### v2

@@ Abstract
Managing data related to a software product and its development poses significant challenges for software projects and agile development teams. Challenges include integrating data from diverse sources and ensuring data quality in light of continuous change and adaptation. To this end, we aimed to systematically explore data management challenges and potential solutions in agile projects. We employed a mixed-methods approach, utilizing a systematic literature review (SLR) to understand the state-of-research followed by a survey with practitioners to reflect on the state-of-practice. In the SLR, we reviewed 45 studies in which we identified and categorized data management aspects and the associated challenges and solutions. In the practitioner survey, we captured practical experiences and solutions from 32 industry experts to complement the findings from the SLR. Our findings reveal major data management challenges reported in both the SLR and practitioner survey, such as managing data integration processes, capturing diverse data, automating data collection, and meeting real-time analysis requirements. Based on our findings, we present implications for practitioners and researchers, which include the necessity of developing clear data management policies, training on data management tools, and adopting new data management strategies that enhance agility, improve product quality, and facilitate better project outcomes.
Keywords:
Data Management Data Quality Agile Software Development
@@ 1 Introduction
Agile software development approaches prioritize flexibility, rapid iteration, and constant adaptation to change. Its dynamic nature can lead to issues related to how data (product, process, project and operational data) is managed in a project, such as integrating diverse data sources and maintaining data quality despite frequent updates and changes franch2019quality . When managing data, agile projects often face problems with data collection and inconsistencies. For example, Fabijan et al. fabijan2016lack observed that data collected by different agile teams often remained isolated in large software-intensive organizations, leading to inefficiencies and repeated efforts, as important feedback was not systematically shared across the organization. A product owner mentioned “it is all in my head more or less”, highlighting the issue of unshared, valuable data. fabijan2016lack . Furthermore, the dynamic nature of privacy requirements and the frequent updates to user stories exacerbate the challenge, as teams often struggle to keep documentation current and comprehensive. Managing data in agile projects poses significant challenges that must be effectively handled to optimize organizational performance graetsch2023dealing . This research investigates data management challenges and solutions in agile software development. We employed a mixed-methods approach using a Systematic Literature Review (SLR) and a survey with practitioners to explore data management challenges and solutions in agile software development.
@@
The remaining sections of this paper are structured as follows: Section 2 provides the background on data management and agile methods. We present the SLR in Section 3 followed by the practitioner survey in Section 4. We discuss the study results, present a list of implications and recommendations in Section 5 followed by a discussion of threats to validity in Section 6. Finally, the conclusions are provided in Section 7.
@@ 2.1 Agile Software Development and Importance of Data Management
Agile software development aims to make incremental progress by continuously incorporating customer feedback and adapting to change through a flexible and collaborative approach. Agile development encourages a development approach focused on the customer’s needs and appreciates the contributions of skilled and empowered teams. Agile methods prioritize people, working software, customer collaboration, and adaptability over rigid processes and planning, as stated in the agile manifesto beck2001manifesto . Agile practices have increased the effectiveness and responsiveness of software development to market demands lee2010toward .
@@
Organizational success is significantly impacted by managing data, which is a crucial resource that helps with decision-making, provides insightful information about customer behaviour, and improves operational efficiency matthies2019towards . Many forms of data (e.g., business or product data) play important roles in agile development, each impacting the development process and the project’s overall success.
@@ 2.2 Types of Data in Agile Software Development
Business data: This data type contains details relevant to business decisions, such as information for forecasting sales, anticipating the need for raw materials, and analyzing consumer behaviour matthies2019towards . This data type enhances productivity and operational efficiency by facilitating direct feedback from users and enabling quick communication between developers and customers in an agile process. This data facilitates iterative cycles of decision-making in agile processes. Teams can use it to prioritize features according to market trends and customer requirements, ensuring that development efforts align with company objectives fabijan2016lack .
@@
Product Data: This data type contains details about the software product itself, including design documents and source code. It measures the software product’s size, complexity, and design structure—all of which significantly impact software quality rahman2013and ; li1999software ; colakoglu2021software . This data provides metrics that can be used to measure code quality, complexity, and technical debt to inform continuous integration and agile delivery processes. Such metrics are critical to sustaining high levels of software quality and enabling quick releases rosenkranz2017supporting .
@@
Process Data: Process data is about the software development process and includes several elements that could affect how much work is needed to build a software system. Understanding and enhancing the software development lifecycle need the use of process data rahman2013and ; li1999software . Process data is essential for continual improvement, which agile techniques emphasize. Teams can modify their procedures for improved results by using the insights it offers about process bottlenecks, team performance, and collaboration efficiency matthies2019towards .
@@
Project Data: This data type focuses on a software project’s general health and state, including resources, risks, budgets, and schedules. Project managers frequently use this data to track the status of their work and inform choices about planning and controlling their projects colakoglu2021software . Agile project management depends on having real-time project data. It assists in monitoring development, controlling risks, and making sure the project stays within budget and on schedule. This data demonstrates the agile principles of transparency and ongoing feedback matthies2019towards .
@@
Operational data is data about a company’s daily activities. It entails privacy engineering, managing personal data, and deploying and running software in cloud environments grunewald2021cloud . Operational data comprises sensitive corporate information, classified documents, and data kept in databases necessary for enterprise operations. Such data is crucial to the day-to-day functioning of the business min2010practices . Within the agile framework, operational data facilitates the daily management of projects.
@@ 2.3 Related Work
To the best of our knowledge, our study is the first review that specifically investigates data management challenges and solutions in agile software projects. An SLR systematically gathers and examines evidence from multiple studies, creating a comprehensive understanding of data management challenges and solutions in agile software development. Previous studies have mostly focused on separating a set of challenges or solutions, which can be systems-specific. For example, Graetsch et al. graetsch2023dealing discussed dealing with data challenges when delivering data-intensive software solutions but did not discuss specific solutions that have been implemented. Similarly, the solution proposed by Rosenkranz et al. rosenkranz2017supporting addresses specific data integration (e.g., harmonizing data from different sources) and data quality challenges (e.g., ensuring accuracy and consistency), highlighting the interconnected nature of these two aspects in effective data management. Another important solution that has not been covered in the previous studies is the use of methods for collecting diverse data through user-centred design pater2018advancing .
@@ 3.1.1 Search Process
We used Scopus to search for relevant studies due to its comprehensive indexing of major publications in software engineering, as highlighted in studies by Mourão et al. and Carrera-Rivera et al. mourao2020performance ; carrera2022conduct . The extensive coverage of Scopus provides access to high-quality, peer-reviewed literature, which is crucial for an in-depth exploration of agile software development. Moreover, data management is a topic at the intersection of multiple disciplines (including software engineering, project management, etc.). Scopus provides more extensive coverage of literature on our study topic as it covers popular publishing venues, including the four major software engineering publication libraries: IEEE Xplore, ACM Digital Library, SpringerLink and ScienceDirect. The search covered literature starts from December 2003 (the date of the first study returned by our search string, as there was no start date limit) until October 2023, when we began consolidating the findings to guide the survey.
@@
These 181 studies proceeded to the next stage of review, as illustrated in Figure 1. Additionally, we conducted a full pilot review process on a sample of these studies. We applied the inclusion and exclusion criteria, performed quality assessments (see Section 3.1.5) and extracted the data to ensure the final selection aligned with our research objectives.
In the search string, we aim to cover all two aspects of the study i.e., agile software development and data management, with the goal of capturing studies that focus on data management (aspects, issues, challenges and solutions) in agile software development. Our final search string (which covered the aspects above) is shown below:
@@
Figure 1: The Review Process
@@ 3.1.2 Inclusion and Exclusion Criteria
We selected studies that met all of the following inclusion criteria:
@@
    Studies that discuss at least one aspect of data management in the context of agile software development (including challenges and solutions).
@@
    Studies that are available in full text for comprehensive analysis.
@@
    Studies that were published in English.
We applied the following exclusion criteria and excluded studies when they met one of these criteria:
  1. 1.
@@
  4. 4.
    Non-research materials such as tutorials, books or workshop summaries.
  5. 5.
@@ 3.1.3 Filtering of Studies
The 181 studies were then screened, where their titles, abstracts, and keywords were examined. Applying the inclusion and exclusion criteria, we filtered out 114 irrelevant studies, leaving 67 relevant studies. To ensure that these studies precisely matched our research questions and scope, we conducted a full-text review, applying the same criteria. This review led to the exclusion of 22 studies from the initial 67. Reasons for exclusion included factors such as language barriers (e.g., one study was written in a non-English language, with only the abstract in English), content type (e.g., one was a tutorial, another an event), or lack of depth or relevance to our scope. Consequently, 45 studies were finally included after a quality assessment (see Section 3.1.5).
@@ 3.1.4 Data Extraction
We utilized a combination of thematic analysis and content analysis to synthesize data from the studies. Thematic analysis was used to identify, analyze, and report themes (data management aspect, challenge, and solution) within the qualitative data extracted from the 45 studies. Initially, a subset of studies (9 studies) were reviewed to identify general data management aspects. This provides a structured understanding of the various aspects of data management within agile software development. Additionally, this foundational knowledge helps identify specific challenges and solutions associated with each aspect, thereby addressing RQs. We categorised the extracted data from the studies (all those items were extracted as free text) as follows:
@@
Subsequently, two authors discussed these aspects in detail to ensure coverage of all challenges and solutions. Relevant data was retrieved and organised ( by careful iterative reading through the complete texts of 45 included studies) for each research question (RQ1: Data management challenges and RQ2: solutions). Challenges were identified based on descriptions of problems faced in managing data within agile projects, as documented in the reviewed studies. Solutions were extracted by noting specific strategies, tools, or methods proposed or implemented to address the identified challenges. For example, the thematic analysis revealed several challenges associated with data integration (see Section 3.2.3), such as data harmonization, semantic heterogeneity, data transformation, and extraction. The identified solutions were also thematically categorized and presented (see Table 2). An iterative process was used to review and validate the extracted data. To guarantee accuracy and consistency, the research team discussed the extracted data from a sample of studies and compared the outcomes. To improve clarity and comprehensiveness, differences were discussed and settled collectively, and the extraction forms were modified as a result. Content analysis was used to classify studies based on the data management aspects they addressed (such as data integration, data quality, data collection, and data analysis). Additionally, we quantified the number of studies discussing each aspect to identify the most and least covered areas. For example, the studies were classified into 15 data management aspects (see Section 3.2.1), listed these aspects, and showed the distribution of studies across these aspects, providing a clear quantification of how many studies addressed each aspect. In addition, the data types discussed in the studies are provided in Table A9. We provide a dataset of all the included studies we have extracted and our detailed analysis in a replication package fawzy_2024_10597818 .

### v4

@@ Abstract
Context: Managing data related to a software product and its development poses significant challenges for software projects and agile development teams. These include integrating data from diverse sources and ensuring data quality amidst continuous change and adaptation.
Objective: The paper systematically explores data management challenges and potential solutions in agile projects, aiming to provide insights into data management challenges and solutions for both researchers and practitioners.
Method: We employed a mixed-methods approach, including a systematic literature review (SLR) to understand the state-of-research followed by a survey with practitioners to reflect on the state-of-practice. The SLR reviewed 45 studies, identifying and categorizing data management aspects along with their associated challenges and solutions. The practitioner survey captured practical experiences and solutions from 32 industry practitioners who were significantly involved in data management to complement the findings from the SLR.
Results: Our findings identified major data management challenges in practice, such as managing data integration processes, capturing diverse data, automating data collection, and meeting real-time analysis requirements. To address these challenges, solutions such as automation tools, decentralized data management practices, and ontology-based approaches have been identified. These solutions enhance data integration, improve data quality, and enable real-time decision-making by providing flexible frameworks tailored to agile project needs.
Conclusion: The study pinpointed significant challenges and actionable solutions in data management for agile software development. Our findings provide practical implications for practitioners and researchers, emphasizing the development of effective data management practices and tools to address those challenges and improve project success.
@@ 1 Introduction
Agile software development approaches prioritize flexibility, rapid iteration, and constant adaptation to change. Its dynamic nature can lead to issues related to how data (product, process, project and operational data) is managed in a project, such as integrating diverse data sources and maintaining data quality despite frequent updates and changes franch2019quality . When managing data, agile projects often face problems with data collection and inconsistencies. For example, Fabijan et al. fabijan2016lack observed that data collected by different agile teams often remained isolated in large software-intensive organizations, leading to inefficiencies and repeated efforts, as important feedback was not systematically shared across the organization. A product owner mentioned “it is all in my head more or less”, highlighting the issue of unshared, valuable data fabijan2016lack . Furthermore, the dynamic nature of privacy requirements and the frequent updates to user stories exacerbate the challenge, as teams often struggle to keep documentation current and comprehensive. Managing data in agile projects poses significant challenges that must be effectively handled to optimize organizational performance graetsch2023dealing . This research investigates data management challenges and solutions in agile software development. We employed a mixed-methods approach using a Systematic Literature Review (SLR) and a survey with practitioners to explore data management challenges and solutions in agile software development.
@@
The remaining sections of this paper are structured as follows: Section 2 provides the background on data management and agile methods. We present the SLR in Section 3 followed by the practitioner survey in Section 4. We discuss the study results and present a list of implications and recommendations in Section 5 followed by a discussion of threats to validity in Section 6. Finally, the conclusion of this study is presented in Section 7.
@@ 2.1 Agile Software Development and Importance of Data Management
Agile software development aims to make incremental progress by continuously incorporating customer feedback and adapting to change through a flexible and collaborative approach. Agile development encourages a development approach focused on the customer’s needs and appreciates the contributions of skilled and empowered teams. Agile methods prioritize people, working software, customer collaboration, and adaptability over rigid processes and planning, as stated in the agile manifesto beck2001manifesto . Agile practices have increased the effectiveness and responsiveness of software development to market demands amajuoyi2024agile .
@@
Organizational success is significantly impacted by managing data, which is a crucial resource that helps with decision-making, provides insightful information about customer behavior, and improves operational efficiency matthies2019towards . Many forms of data (e.g., business or product data) play important roles in agile development, each impacting the development process and the project’s overall success.
@@ 2.2 Types of Data in Agile Software Development
Business data: This data type contains details relevant to business decisions, such as forecasting sales, anticipating the need for raw materials, and analyzing consumer behavior matthies2019towards . This data type enhances productivity and operational efficiency by facilitating direct feedback from users and enabling quick communication between developers and customers in an agile process. This data facilitates iterative cycles of decision-making in agile processes. Teams can use it to prioritize features according to market trends and customer requirements, ensuring that development efforts align with company objectives fabijan2016lack . For example, during the COVID-19 pandemic, population flow data (i.e., aggregated data from mobile phone location tracking or similar systems that show how people move between areas) was used to identify high-risk regions, allowing targeted health measures and better resource planning huang2021leveraging .
@@
Product Data: This data type contains details about the software product itself, including design documents and source code. It measures the software product’s size, complexity, and design structure, which all significantly impact software quality rahman2013and ; li1999software ; colakoglu2021software . This data provides metrics that can be used to measure code quality, complexity, and technical debt to inform continuous integration and agile delivery processes. Such metrics are critical to sustaining high levels of software quality and enabling quick releases rosenkranz2017supporting . For example, privacy requirements can be integrated into user stories to ensure compliance with the data protection law canedo2022guidelines , which helped agile teams assess and implement data anonymization measures and other privacy safeguards during the development process.
@@
Process Data: Process data is about the software development process and includes several elements that could affect how much work is needed to build a software system. Understanding and enhancing the software development lifecycle need the use of process data rahman2013and ; li1999software . Process data is essential for continual improvement, which agile techniques emphasize. Teams can modify their procedures for improved results by using the insights it offers about process bottlenecks, team performance, and collaboration efficiency matthies2019towards . For example, Fabijan et al. fabijan2016lack analyzed task handover times between development and testing phases to identify delays and inefficiencies. This data revealed bottlenecks caused by a lack of data sharing, leading to repetitive work and suboptimal decision-making. Addressing these issues streamlined the development process and improved overall team collaboration.
@@
Project Data: This data type focuses on a software project’s general health and state, including resources, risks, budgets, and schedules. Project managers frequently use this data to track the status of their work and inform choices about planning and controlling their projects colakoglu2021software . Agile project management depends on having real-time project data. It assists in monitoring development, controlling risks, and making sure the project stays within budget and on schedule. This data demonstrates the agile principles of transparency and ongoing feedback matthies2019towards . For example, metrics like the mean time between failures (MTBF) have been used to predict and address potential software errors during agile sprints. Such project data supports resource allocation and task prioritization, improving project outcomes batarseh2018predicting .
@@
Operational data is data about a company’s daily activities. It entails privacy engineering, managing personal data, and deploying and running software in cloud environments grunewald2021cloud . Operational data comprises sensitive corporate information, classified documents, and data kept in databases necessary for enterprise operations. Such data is crucial to the day-to-day functioning of the business min2010practices . Within the agile framework, operational data facilitates the daily management of projects. For example, agile methods were applied to create patient registries using electronic health record data for chronic disease management kannan2017rapid . This operational data, managed by the organization deploying the system, supported real-time patient monitoring and reporting kannan2017rapid . In contrast to operational data, which focuses on supporting daily activities, business data is used for strategic decision-making (e.g., aggregating population flow data during COVID-19 to identify high-risk areas and optimize resource allocation).
@@ 2.3 Related Work
To the best of our knowledge, our study is the first review that specifically investigates data management challenges and solutions in agile software projects. An SLR systematically gathers and examines evidence from multiple studies, creating a comprehensive understanding of data management challenges and solutions in agile software development. Previous studies have mostly focused on separating a set of challenges or solutions, which can be systems-specific. For example, Graetsch et al. graetsch2023dealing discussed dealing with data challenges when delivering data-intensive software solutions but did not discuss specific solutions that have been implemented. Similarly, the solution proposed by Rosenkranz et al. rosenkranz2017supporting addresses specific data integration (e.g., harmonizing data from different sources) and data quality challenges (e.g., ensuring accuracy and consistency), highlighting the interconnected nature of these two aspects in effective data management. Another important solution not covered in the previous studies is the use of methods for collecting diverse data through user-centered design pater2018advancing .
@@ 3.1.1 Search Process
We used Scopus to search for relevant studies due to its comprehensive indexing of major publications in software engineering, as highlighted in studies by Mourão et al. mourao2020performance and Carrera-Rivera et al. carrera2022conduct . The extensive coverage of Scopus provides access to high-quality, peer-reviewed literature, which is crucial for an in-depth exploration of agile software development. Moreover, data management is a topic at the intersection of multiple disciplines (including software engineering, project management, etc.). Scopus provides more extensive coverage of literature on our study topic as it covers popular publishing venues, including the four major software engineering publication libraries: IEEE Xplore, ACM Digital Library, SpringerLink, and ScienceDirect. The search covered literature starts from December 2003 (the date of the first study returned by our search string, as there was no start date limit) until October 2023, when we began consolidating the findings to guide the survey.
@@
These 181 studies proceeded to the next stage of review, as illustrated in Figure 1. Additionally, we conducted a full pilot review process on a sample of these studies. We applied the inclusion and exclusion criteria, performed quality assessments (see Section 3.1.5), and extracted the data to ensure the final selection aligned with our research objectives.
In the search string, we aim to cover all two aspects of the study, i.e., agile software development and data management, to capture studies that focus on data management (aspects, issues, challenges, and solutions) in agile software development. Our final search string (which covers the aspects above) is shown below:
@@
To ensure the completeness and reliability of our study selection, we adopted an informal snowballing approach as a supplementary validation step. Specifically, we examined the references of a subset of the top nine cited studies, treating them as a quasi-gold standard zhang2011identifying , i.e., a set of highly relevant studies. This backward snowballing approach allowed us to cross-check our dataset and confirm that no key studies were missed due to search string limitations. After reviewing these references, we found that no additional studies met our inclusion criteria, reinforcing the robustness and comprehensiveness of our selection process. We conducted this search before extracting the results (see Figure 1).
Figure 1: The Systematic Literature Review Process
@@ 3.1.2 Inclusion and Exclusion Criteria
We selected studies that discuss at least one aspect of data management in the context of agile software development (including challenges and solutions). We applied the following exclusion criteria:
@@
    Studies that are not available in full text (e.g., abstracts only).
@@
    Studies that are not published in English.
@@
  6. 6.
    Non-research materials such as tutorials, books, or workshop summaries.
  7. 7.
@@ 3.1.3 Filtering of Studies
We screened all 181 studies and examined all titles, abstracts, and keywords used. The first author conducted the initial screening, applying the inclusion and exclusion criteria to ensure relevance. The second co-author cross-validated the screening to verify consistency and accuracy. By applying the inclusion and exclusion criteria, we filtered out 114 irrelevant studies, leaving 67 relevant studies. To ensure that these studies precisely matched our research questions and scope, we conducted a full-text review, applying the same criteria. The full-text review followed the same structured approach: the first co-author conducted the initial review, and the second co-author cross-validated the results. This review led to the exclusion of 22 studies from the initial 67. Reasons for exclusion included factors such as language barriers (e.g., one study was written in a non-English language, with only the abstract in English), content type (e.g., one was a tutorial, another an event), or lack of depth or relevance to our scope.
The filtration process was conducted independently by the first co-author in the initial phase, followed by cross-validation by the second co-author. Any disagreements were resolved through discussions and consensus among all four authors, ensuring transparency and rigor. After applying the filtration process, we included a total of 45 studies (see Section 3.1.5).
@@ 3.1.4 Data Extraction
We utilized a combination of thematic analysis and content analysis to synthesize data from the studies. Thematic analysis braun2006using was used to categorize, analyze, and report themes (data management aspect, challenge, and solution) within the qualitative data extracted from the 45 studies. Initially, a subset of studies (9 studies) were reviewed to identify general data management aspects. This provides a structured understanding of the various aspects of data management within agile software development. Additionally, this foundational knowledge helps identify specific challenges and solutions associated with each aspect, thereby addressing RQs. We categorized the extracted data from the studies (all those items were extracted as free text) as follows:
@@
Our data extraction process was carried out collaboratively to ensure accuracy and reliability. The first co-author performed the initial data extraction, while the second co-author cross-validated the extracted data to ensure consistency and alignment with the research objectives. The last two co-authors reviewed the extracted data, offering insights and refinements to enhance the comprehensiveness of the results.
@@
Subsequently, any disagreements in the extracted data were resolved through discussions among all authors until a consensus was reached. Relevant data was retrieved and organized by careful iterative reading through the complete texts of 45 included studies for each research question. Data management challenges (RQ1) were identified based on descriptions of problems faced in managing data within agile projects, as documented in the reviewed studies. The solutions (RQ2) were extracted by noting specific strategies, tools, or methods proposed or implemented to address the identified challenges. For example, thematic analysis revealed several challenges associated with data integration (see Section 3.2.4), such as data harmonization, semantic heterogeneity, data transformation, and extraction. The identified solutions were also thematically categorized and presented (see Table 3). An iterative process was used to review and validate the extracted data. To ensure accuracy and consistency, the two coders discussed the extracted data from a sample of studies and compared the outcomes. They also discussed any differences to settle any disagreements in classification. We also employed content analysis krippendorff2018content to analyze studies based on the data management aspects they addressed. Additionally, we quantified the number of studies discussing each aspect to identify the challenges and solutions that were covered the most and least. For example, the studies were classified into 15 data management aspects (see Section 3.2.1), listed these aspects, and showed the distribution of studies across these aspects. In addition, the data types discussed in the studies are provided in Table A9. We provide a dataset of all the included studies we have extracted and our detailed analysis in a replication package fawzy_2024_10597818 .
Both thematic and content analysis were important to ensure a comprehensive and balanced synthesis of the findings. Thematic analysis allowed us to gain deep insights into the studies by identifying recurring patterns and themes. For instance, it highlighted that challenges with data harmonization often overlap with issues in ensuring data accuracy and consistency due to the lack of standardization in data integration pipelines. On the other hand, content analysis enabled us to quantify the prevalence of these challenges across the studies, revealing their relative significance, such as data integration being addressed in 19 studies (see Section 3.2.3). This mixed-methods approach combined the depth of qualitative insights with the validation of quantitative evidence. Integrating qualitative and quantitative methods enables a deeper understanding of complex phenomena by combining the detailed exploration of qualitative data (e.g., uncovering relationships between challenges) with the ability of quantitative data to validate findings across a broader range of studies (e.g., identifying widely addressed challenges) fetters2013achieving . The prioritization of data management challenges (see Section 3.2.3) is an example of the complementary contributions of both methods. The following are the combined thematic and content analysis steps that we follow:
  • Understanding the Data: We first familiarized ourselves with the data by reading over the studies included in the SLR after applying the filtration process. . At this stage, content analysis focused on identifying broad categories or topics within the filtered dataset noted for their frequency and relevance to inform the next steps in analysis.
  • Generating Initial Codes: We have iteratively read each of the included studies to identify the possible codes and themes. The first author developed the initial codes, which the second author then reviewed for consistency. These initial codes highlighted key patterns in the data and formed the basis for further analysis. We employed content analysis to quantify mentions of key terms or challenges.
  • Searching for Themes: Related codes were grouped into broader themes. For instance, “semantic heterogeneity” and “data harmonization” were grouped under the theme “Data Integration”. Thematic grouping was supported by calculating the prevalence of codes. For example, the theme “Data Integration” emerged as the most frequently discussed aspect, appearing in 19 studies. This quantification ensured that identified themes were representative of the dataset.
  • Reviewing Themes: The relationships between themes were examined. For example, challenges with data integration often overlapped with data quality issues, reflecting their interconnected nature in agile projects. The data was cross-checked against the initial studies to ensure accuracy and consistency. The extracted quantitative data from content analysis validated thematic findings by showing overlap in study coverage. For example, several studies discussed both ”data integration“ and ”data quality“, underscoring the interconnected nature of these challenges. This helped ensure the themes were comprehensive and grounded in evidence.
  • Defining and Refining Themes: Each theme was clearly defined and named. For example, “Data Integration” was described as the process of combining data from various sources to form a unified dataset.Content analysis helped quantify the prevalence of the challenges across the studies. For instance, the “Data Integration” theme was supported by 4 studies discussing “data harmonization and interoperability” and 4 studies addressing “semantic heterogeneity”, as shown in Table 3.2.3. Similarly, the “Data Quality” theme was supported by 5 studies focusing on “ensuring data accuracy and consistency”.